OpenStudy (anonymous):

Fine grains of beach sand are assumed to be spheres of radius 38.9 µm. These grains are made of silicon dioxide which has a density of 2600 kg/m3 . What is the mass of each grain of sand? Answer in units of kg.

OpenStudy (anonymous):

well mass = density * volume and the volume of a sphere is \[Volume = (4/3)*\Pi*Radius ^{3}\]

OpenStudy (anonymous):

so \[Volume = (4/3)*\Pi*(38.9*10^{-6})^{3}\]\[Volume = 0.2465683979*10^{-12} m ^{3}\]finally\[mass = 2600*0.2465683979*10^{-12} =0.641kg\]

OpenStudy (anonymous):

\[6.41∗10^{−10}kg\] your last calculation is off by a factor of a billion

OpenStudy (anonymous):

6.41*1/10^10

OpenStudy (anonymous):

Consider a cube whose sides are 1.18 m long. How many kg of sand would it take for the total surface area of all the grains of sand to equal the surface area of the cube?